OpenStudy (anonymous):

Factor the polynomial completely. (If it is not factorable using integers, enter NF.) 10 + 3x − x^2

OpenStudy (anonymous):

\[10 + 3x - x^{2} = -\left( x^{2}-3x-10 \right)=-\left( x-5 \right)\left( x+2 \right)\]

OpenStudy (anonymous):

Thanks, the way they set the problem up confused me.